Its been a while since I have done functions, so I'm a little confused on this one.
if f(x)=4x/1-x and g(x)=2/x then f(g(x))=
the answer is supposed to be 8/x-2 but i got 8/1-x
I'm also a bit confused on this one as well
sqrt(a)/1+sqrt(a)=
the answer is supposed to be sqrt(a)-a/1-a but I just got 1

Answer by rothauserc(4718)

1)if f(x)=4x/1-x and g(x)=2/x then
f(g(x))= f(2/x) = 4*(2/x) / (1 - 2/x)
= (8/x) / ((x-2) / x)
= (8/x) * x / (x-2)
= 8 / (x - 2)
2) sqrt(a)/1+sqrt(a)= a^(1/2) / (1 + a^(1/2))
now multiply the denominator and numerator by (1 - a^(1/2)) and we get
= (a^(1/2) * (1-a^(1/2)) / (1 + a^(1/2)) * (1 - a^(1/2))
= (a^(1/2) - a) / (1 - a)